OpenStudy (anonymous):

v=1/3bh solve the formula for the specified variable? I said h=3v/b, is this correct?

OpenStudy (anonymous):

solving for h we should get h=3v/b so yes you are right ;)

OpenStudy (anonymous):

v=(1/3)*b*h, 3v=bh, 3v/b=h There you go.

jhonyy9 (jhonyy9):

yes this is right